OpenStudy (anonymous):

pls need help on hw 12??????

OpenStudy (anonymous):

HWP3 - first 5 filter: 10Hz and 19kHz. signals around 1kHz, to be amplified by about a factor of 2. 0.68μF and a 390pF capacitor. resistors between 100Ω and 100kΩ. RS=20kΩ. 1) First find C1=0.68*10^-6 = 6.8e-7 (answer of part 4) Then R1*C1=1/(2*pi*f) R1*C1=1/(2*pi*10) so R1= 23417 2)First find C2=390*10^-12 = 3.9e-10 (answer of part 5) R2*C2=1/(2*pi*19*10^3) R2= 21489 3) Rs = Rf = 20kΩ = 2000 Ω

OpenStudy (anonymous):

Remaining answers HWP3 part 6 : R3= R1 part 7 : R4= R2 part 8: c3=c1 part 9 : c4= c2 part 9 :Midband gain = -­(R4/R3)

OpenStudy (anonymous):

some one asked me the value of pi. its always 3.14 buddy!